# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ III.2 Transport in trapped ion chains
The possibility to perform normal mode spectroscopy of an ion chain and to address and monitor a single ion within the chain make these systems a powerful laboratory for studying thermodynamics of and heat transport in many-body systems at the quantum level. This has led to investigations of coherent transport of excitations along ion chains varying from few to dozens of ions, and temperatures ranging from the Doppler cooling limit to near zero point energies.
Figure 41: Schematic of the energy transport experiment of 261. (a) An ion at the chain edge is excited by a pulsed beam (blue). Following free evolution, the energy is read out along the chain with a probe beam (red). (b) CCD image of the ion chain with 37 ions. The ions are 40Ca+ and initially Doppler cooled. The pseudopotential frequencies of the Paul trap are $(\omega_{x},\omega_{y},\omega_{z})/(2\pi)=(2.25,2.0,0.153)$ MHz. The vibrations are locally excited along the $x$ direction, corresponding to the steepest confining potential. From 261.
Figure 42: Energy transport in a chain of five ions. The leftmost ion is given a kick and its energy is measured after a subsequent evolution time $\tau$ as shown in (a). Energy revivals approaching the initial energy are observed. Panel (b) shows the energy of the rightmost ion: initially unexcited, the energy from the kick is rapidly transferred across the chain. Panels (c) and (d) show the energies of the ions from molecular dynamics simulations. From 261.
Coherent transport of vibrational excitations was studied in ion chains in linear Paul traps by [139, 261]. We discuss in more details the experiment of [261], as this is representative of the state of the art reached on the observation of these dynamics. Here, transport of a transverse vibrational excitation was measured along chains of 40Ca+ in a linear Paul trap with the number of ions ranging from 5 to 37, as illustrated in Fig. 41. After initializing the chain motion to Doppler cooling temperatures, an edge ion was excited by a laser pulse with a finite duration. The intensity of the pulse was modulated at the trap radial frequency, thereby locally exciting a transverse vibration. The dynamics of energy propagation was monitored by resolving individual ions along the chain by driving resonant laser pulses and extracted from measurement of the time-evolution of the probability that the ion is in the electronic ground state. The measured data is consistent with a model of coupled oscillators (or equivalently, with a harmonic crystal). In a quantized form, the model for the vibrations dynamics along the transverse $x$-direction reads:
@@
where $i$ labels the ions along the chain ($i=1,\ldots,N$), and $a_{i}$ is the boson annihilation operator of a quantum of transverse vibration at site $i$. In this representation, an excitation hops between sites with a non-local hopping amplitude $t_{ij}$. The frequency of the local oscillators is the transverse trap frequency $\omega_{x}$, here along the $x$-direction, and includes a position-dependent frequency shift due to the Coulomb interactions: $\omega_{x,i}=\omega_{x}-\sum_{i\neq j}t_{ij}$. In the experiment, the hopping element between the leftmost ion and its neighbor ranges in the interval $t_{ij}\in 2\pi\times(2.21,6.7)$ KHz, where the smallest (largest) value is for a chain of 5 ions (25 ions). The measurements show that the time required for the excitation to propagate across the entire chain is comparable with the inverse of the coupling rate between nearest neighbors. This behavior was interpreted as a manifestation of the long-range hopping, which permits the excitation to directly hop to a distant ion and favors transfer to the other edge ion, since it has the same local frequency and the coupling is resonant. These features also explain why adjacent ions can get very different energy excitations.
@@
Revivals of the excitation energy, visible in Fig. 42, were reported even for long chains. The revivals persisted for times of the order of 40 ms with measured time scales which seem weakly dependent on the number of ions. Their decay was attributed to the change of the radial trap frequency over this time. These revivals can be understood as due to the rephasing of multiple phonon normal modes and signal the coherent dynamics of the propagating excitation, showing that the chain had not thermalized. In these settings, in Ref. [1] the authors measured the spreading and refocussing of a vibrational excitation on a chain composed by 42 ions, by means of a Ramsey-type of measurement as in [69]. They could access the autocorrelation functions of the phonons as well as the quantum discord. The experiments so far described were performed with cold, yet thermal, chains. Deep in the quantum regime, transport of a single quantum of vibration was first realized in Ref. [307] on a chain composed of four 40Ca+. The experiment reported coherence time exceeding 10 ms. The experimental measurements are in excellent agreement with the predictions of Eq. (66).
@@
Heat transport in ion chains is the subject of several theoretical studies. In Ref. [194] the energy distribution along the chain was determined when two ions in the chain realize point contacts to heat baths at different temperatures. This configuration can be implemented by continuously laser cooling the two individual ions to different stationary temperatures. A rich behavior was predicted as a function of the cooling rate and of the position of the laser-cooled ions within the chain. Heat transport through an ion chain that undergoes the linear to zigzag transition was experimentally measured in Ref. [261]. A theoretical study reported that the onset of zigzag order gives rise to a qualitative change in transport, from ballistic to diffusive [270]. The origin of the diffusive behavior was attributed to the nonlinearity induced by coupling between the longitudinal and transverse phonons to the zigzag modes. They also found that the heat conductivity is maximized near the critical point, where the transverse phonon becomes gapless. Transport in a chain embedding two mass impurity defects was theoretically analyzed in Ref. [111, 306]. Here, it is argued that for certain mass ratios, the motion of the mass defects can become entangled via the resonant exchange of vibrational excitations with the rest of the chain. Heat transport in the presence of a topological defect was investigated in [316] and is reviewed in Sec. III.3. The properties of transport using the radial modes of a linear chain are at the basis of proposals for quantum information with trapped ion chains [288].
@@
For sake of completeness, we mention transport experiments on the coherent exchange of phononic excitations realized in microtraps, where the ion structure is imposed by the external potential [38, 132, 131]. This experimental progress demonstrates the capability of controlling and designing phonon hopping between the ions and sets the basis for simulating many-body dynamics via the phonons such as the Bose-Hubbard model [253] as well as probing quantum features of transport with bosons [21]. Further work has been devoted to transport controlled by the coupling with the internal degrees of freedom of the ions. We finally mention the realization of vibrationally-assisted transfer of energy [126] and the simulation of an effective Jaynes-Cummings type of dynamics [286] in order to achieve effects such as phonon-blockade and designed phonon-phonon interactions [71, 237].
@@ III.3 Topological defects and metastability
Figure 43: Experimentally observed kinks in an ion chain of about 30 ions. The kinks (red ellipses) are created during a radial quench of a crystal and imaged with exposure times of 40 ms. (a) Uniform zigzag and localized kink for trap aspect ratio $\nu/\nu_{t}\approx 8$. (b) Zigzag and extended kink for $\nu/\nu_{t}\approx 5.5$. Adapted from 243.
Manifestations of prethermalization can be observed in ion chains in the form of dislocations. Figure 43 reports some examples of dislocations (defects) on a plane in a zigzag chain imaged on a CCD camera. It does not show defects composed of dislocation in three dimensions, see [211, 234].
Defects are typically created by means of temporal transformations of the trap potential, as we will review in the next section. They can also be generated by the laser cooling dynamics. In [211] the formation of defects was observed by laser cooling an ion cloud of 24Mg+ ions in a linear Paul trap to the millikelvin regime. In this regime, for the chosen trap aspect ratio, the ordered phase was a zigzag configuration. By imaging the ions positions via resonance fluorescence collected with a CCD camera, zigzag configurations with defects were measured at the center of the trap with a rate of occurrence depending on the number of ions. These configurations decay into the regular zigzag chain by propagation of the defect from the center to the edges of the zigzag chain after a time of the order of 10 s. This time scale shall be compared with the inverse of the axial trap frequency $1/\nu$ and exceeds it by 5 orders of magnitude.
@@
This variety of defects and dynamics was also reported in experiments where the trap aspect ratio is slowly varied across the linear-zigzag transition while the chain is continuously laser cooled [258, 320, 243, 103]. They are reproduced by molecular dynamics simulations where the trap was described in the pseudopotential approximation, showing that micromotion is not essential in determining this behavior. Their metastable character identifies them as topological defects. Comparison between molecular dynamics and experimental data lead to the identification of kink-kink interactions in chains containing two or more kinks [185].
@@
The kink classification was systematically performed in [185] using tools of topological degree theory. Figure 44 illustrates the metastable (solid line) and unstable (dashed line) equilibrium configurations of one planar kink, which were theoretically calculated for a chain of 31 ions as a function of the trap aspect ratio [185]. The stability is determined by the analysis of the eigenvalues of the corresponding Hessian matrix; a bifurcation occurs everytime an eigenvalue of the Hessian changes sign.
@@
The kink’s dynamics is characterized by localized modes, which are typically constituted by the translational mode and by internal modes. Theoretical analysis of the internal modes for the extended and the localized kinks predicted well resolved spectroscopic features and long coherence times also at Doppler cooling temperatures [183]. The translational energy of the topological defects were studied by calculating the Peierls-Nabarro potential, namely, the energy of the soliton as a function of the location of the kink centre [243, 185]. These studies showed how the presence of mass impurities gives rise to a further class of topological defects [185] and that, combined with static electric fields, they can allow for deterministic manipulation and creation of kinks. Structural phase transitions in the presence of mass impurities have been characterised in [269].
@@
Figure 44: Schematic illustration of bifurcations and stability of one-kink configurations in a planar crystal of 31 ions. The possible ion configurations are depicted as a function of the ratio of confinement in the axial and one of the transverse directions, where confinement in the second transverse direction is assumed to be very strong. Solid lines indicate stable configurations (local minima of the potential), dashed lines are unstable solutions. The arrows illustrate the spatial oscillations of the normal mode which crosses zero energy at the bifurcation and generates the bifurcating solutions. From 185.
The spectroscopic features of the localized soliton were observed in [40] by parametric heating. The measurements of the kink motion suggest directed transport as a function of the topological charge. The stability of the extended kink as a function of the trap ratio and its dynamics were extensively analysed in [173], unveiling features of the sliding to pinning transition, see Sec. III.5.
The possibility to control topological defects in ion chains have inspired several theoretical proposals. In Ref. [184] discrete solitons generate entanglement between the ions in a ring. Numerical studies of transport of energy in ion chains with a topological defect reported energy localization at the kink [315]. When the edge ions are heat point contacts at two different temperature, the kink can substantially modify the heat flow giving rise to a relatively steep temperature gradient at its location [314]. These investigations also indicate that the interactions between kink and normal modes of the chain affect the transport properties.
Beyond linear trap geometries, numerical studies reported topological solitons [48] in tubular structures formed by ions in multipole traps [52]. Dislocations have been reported in 3 dimensional crystals in Penning traps [216]. An extensive characterization of defects and dislocations in planar and higher dimensional ion Coulomb crystals is to date an open problem.
Chains containing mass or charge impurities are routinely realised in the ion-trap laboratories, see for instance [107, 168, 137]. The deterministic preparation of dual species ion crystals has been demonstrated in [338]. By positioning the impurities within the chain the dispersion spectrum can be designed [300], finding application for quantum logic [226, 144] and metrology [283, 144].
@@ III.4 Slow quenches and Kibble-Zurek mechanism
Experimental imaging of dislocations in ion Coulomb crystals opens the possibility to verify the so-called Kibble-Zurek (KZ) hypothesis [170, 340, 47]. This hypothesis was first formulated in an attempt to describe inhomogeneities at the cosmological scale as defects formed during the slow expansion of the universe [170] and then proposed as a paradigm for adiabatic expansion across a continuous, symmetry breaking phase transition [340]. The KZ hypothesis predicts the scaling of the excess heat produced by a slow ramp of a control field across a second–order phase transition, and connects herewith the equilibrium critical exponents with features of the out-of-equilibrium dynamics.
It has been argued that the symmetry-breaking linear-zigzag transition is an ideal testbed for the Kibble-Zurek hypothesis [46, 70]. The adiabatic expansion can be realized by performing slow ramps of the trap aspect ratio across the structural transition. Excess heat can be determined through the density of topological defects created by crossing the critical point. Experiments following this protocol reported a scaling of the excess heat with the ramp velocity that appears to follow power-law behavior [258, 320, 103]. In this Section we review the theoretical hypothesis and the experimental results.
The model below is adapted for a linear trap where the chain is along the $x$-axis and the zigzag chain is in the xyx-y plane, while the motion along $z$ is frozen out. It is taken from [69]. We refer the reader also to a related semiclassical model by [257]. The dynamics of defect formation is modelled by a Langevin equation for the transverse displacement field $\psi(x,t)$:
@@
This equation assumes that the field dynamics is determined by the Landau potential Eq. (45) (with $V_{0}$ given by Eq. (43)) and includes the dissipative dynamics of laser cooling, with damping at rate $\eta$ and the corresponding Langevin force $\varepsilon(t)$ (recall: $\langle\varepsilon(t)\rangle=0$ and $\langle\varepsilon(t)\varepsilon(t^{\prime})\rangle=(2\eta k_{\mathrm{B}}T/m)\delta(t-t^{\prime})$ where $T$ is the stationary temperature of Doppler cooling [302, 222]). The first two terms on the left-hand side describe a wave-like propagation with velocity $v(x)$, which is position dependent when the chain’s charge distribution is inhomogeneous (see e.g. Eq. (38)). For a sufficiently long chain one can define a position-dependent critical trap frequency $\nu_{t}^{(c)}(x)$, at which the chain becomes locally unstable: In a linear Paul trap, at the center the functional is maximum and takes the value $\nu_{t}^{(c)}(0)=\nu_{t}^{(c)}$ (see Sec. II.1.4), while it decreases with the distance $|x|$ from the center. The parameter $\delta(x,t)\approx\nu_{t}(t)^{2}-\nu_{t}^{(c)}(x)^{2}$ gives the amplitude of the local force in the harmonic approximation while term $2\mathcal{A}\psi^{3}$ is the force in next order, see Eq. (43).
Figure 45: Density of defects $d$ for an ion chain in a linear trap as a function of the inverse of the sweeping rate (a) in the overdamped and (b) in the underdamped regime. The circles are obtained from a molecular dynamics simulation with $50$ ions in a linear Paul trap, the defects are only counted in the central 30 ions. The red line is the fitting function $d\sim\tau_{Q}^{-\alpha}$ giving (a) $\alpha=0.995$ and (b) $\alpha=1.427$. Both results are consistent with the predictions of the inhomogeneous KZ mechanism. The curves saturate mostly because of finite size effects. From 46.
@@
For slower quenches, spatial inhomogeneities along the chain, such as the non-uniform distance between the ions in a linear Paul trap, become relevant. The speed of the propagating front, at which the critical value $\nu_{t}^{(c)}(x)$ is crossed along the chain, shall be now compared with the velocity $v(x)$, at which a perturbation propagates from the center to the edges of the chain. Clearly, large front velocities will give rise to nucleation of defects along the chain while instead large sound velocities $v(x)$ will causally connect the different regions and tend to suppress them. Both quantities depend on the local density but with different functional behaviors. This leads to different scaling exponents, namely, $\alpha=1$ for the overdamped regime and $\alpha=4/3$ for the underdamped dynamics [46, 70]. These scalings are recovered in molecular dynamics simulations, as visible in Fig. 45.
@@
Figure 46: Defect formation rates in a chain of 16 40Ca+ ions by changing the axial trap frequency across the linear-zigzag transition. Double-logarithmic plot of the measured density of defects $d$ versus the rate of change of $\omega_{\rm ax}$ at the critical point, where $\omega_{\rm ax}$ is the trap confinement frequency along the direction of the ion chain. All circles correspond to 60000 measurements at a radial trap anisotropy of $1.03$. The fitted function of the form $d\propto 1/\tau_{Q}^{\beta}$ (red line) gives an exponent $\beta=2.68\pm 0.06$. The constant offset visible at lower ramping rates stems from background gas collisions. The saturation is due to the maximum number of defects being limited by the system size. For comparison, the rate of defects measured with a higher radial trap anisotropy of 1.05 is plotted (squares), showing a loss of defects but a similar fitted exponent of $\beta=2.62\pm 0.15$ (dotted line). Solid data points are used for the fits. The blue stars depict the result of molecular dynamics simulations. Adapted from 320
@@
Experimentally the slow quench was performed in three different platforms, with different number of ions and ion species. The trap aspect ratio was varied across the linear-zigzag transition by either sweeping the transverse [258] or the axial frequency [320], or both simultaneously [103]. In all experiments the ions were initially cooled to the Doppler limit and the cooling laser was on during the whole duration of the quench. The density of defects was measured after the quench using a CCD camera per sweep and averaging over several sweeps reporting consistent results. Figure 46 shows the statistics of the kinks experimentally detected in [320]. The solid and dashed lines show the function $d\sim\tau_{Q}^{-8/3}$, which fits the interval in which the sweep rate exceeds the rate of background collisions but is still sufficiently small so that defect generation does not saturate. It has also been reported in the defect statistics measured in [258] on a larger parameter interval, where the power law scaling is also reproduced by accurate molecular dynamics simulations. Furthermore, it is compatible with the measurement performed in [103]. This power-law scaling has been attributed to a form of KZ mechanism, even though a theoretical model justifying this scaling has not been yet identified, see [47].

### v2

@@ III.2 Transport in trapped ion chains
The possibility to perform normal mode spectroscopy of an ion chain and to address and monitor a single ion within the chain make these systems a powerful laboratory for studying thermodynamics of and heat transport in many-body systems at the quantum level. This has led to investigations of coherent transport of excitations along ion chains varying from few to dozens of ions, and with temperatures ranging from the Doppler cooling limit to near zero point energies.
Figure 41: (Color online) Schematic of the energy transport experiment of 292. (a) An ion at the chain edge is excited by a pulsed beam (horizontal blue arrow). Following free evolution, the energy is read out along the chain with a probe beam (vertical red arrow). (b) CCD image of the ion chain with 37 ions. The ions are 40Ca+ and initially Doppler cooled. The pseudopotential frequencies of the Paul trap are $(\omega_{x},\omega_{y},\omega_{z})/(2\pi)=(2.25,2.0,0.153)$ MHz. The vibrations are locally excited along the $x$ direction, corresponding to the steepest confining potential. From 292.
Figure 42: Energy transport in a chain of five ions. The leftmost ion is given a kick and its energy is measured after a subsequent evolution time $\tau$ as shown in (a). Energy revivals approaching the initial energy are observed. Panel (b) shows the energy of the rightmost ion: initially unexcited, the energy from the kick is rapidly transferred across the chain. Panels (c) and (d) show the energies of the ions obtained from molecular dynamics simulations. From 292.
Coherent transport of vibrational excitations was studied in ion chains in linear Paul traps by [151, 292]. We discuss in more details the experiment of [292], as this is representative of the state of the art reached on the observation of these dynamics. Here, transport of a transverse vibrational excitation was measured along chains of 40Ca+ in a linear Paul trap with the number of ions ranging from 5 to 37, as illustrated in Fig. 41. After initializing the chain motion to Doppler cooling temperatures, an edge ion was excited by a laser pulse with a finite duration. The intensity of the pulse was modulated at the trap radial frequency, thereby locally exciting a transverse vibration. The dynamics of energy propagation was monitored by resolving individual ions along the chain using focused, near resonant laser pulses to convert motional excitation to electronic excitation, which is routinely measured. The measured data is consistent with a model of coupled oscillators (or equivalently, with a harmonic crystal). In a quantized form, the model for the vibrational dynamics along the transverse $x$-direction reads:
@@
where $i$ labels the ions along the chain ($i=1,\ldots,N$), and $a_{i}$ is the boson annihilation operator of a quantum of transverse vibration at site $i$. In this representation, an excitation hops between sites with a non-local hopping amplitude $t_{ij}$. The frequency of the local oscillators is the transverse trap frequency $\omega_{x}$, here along the $x$-direction, and includes a position-dependent frequency shift due to the Coulomb interactions: $\omega_{x,i}=\omega_{x}-\sum_{i\neq j}t_{ij}$. In the experiment, the hopping element between the leftmost ion and its neighbor ranges in the interval $t_{ij}\in 2\pi\times(2.21,6.7)$ kHz, where the smallest (largest) value is for a chain of 5 ions (25 ions). The measurements show that the time required for the excitation to propagate across the entire chain is comparable with the inverse of the coupling rate between nearest neighbors. This behavior was interpreted as a manifestation of the long-range hopping, which permits the excitation to directly reach to a distant ion. In this configuration, long-range hopping favors direct transfer between the edge ions, since they possess the same local frequency. These features also explain why adjacent ions could get very different energy excitations.
@@
Revivals of the excitation energy are visible in Fig. 42. They were reported even for long chains. The revivals persisted for times of the order of 40 ms with measured time scales which seem weakly dependent on the number of ions. Their decay was attributed to the change of the radial trap frequency over this time. These revivals can be understood as due to the rephasing of multiple phonon normal modes and signal the coherent dynamics of the propagating excitation, showing that the chain had not thermalized. In these settings, in Ref. [1] the authors measured the spreading and refocussing of a vibrational excitation on a chain composed of 42 ions, by means of a Ramsey-type of measurement as in [74]. They could access the autocorrelation functions of the phonons as well as the quantum discord. The experiments so far described were performed with cold, yet thermal, chains. Deep in the quantum regime, transport of a single quantum of vibration was first realized in Ref. [342] on a chain composed of four 40Ca+. The experiment reported coherence time exceeding 10 ms. The experimental measurements were in excellent agreement with the predictions of Eq. (66).
@@
Heat transport in ion chains is the subject of several theoretical studies. In Ref. [213] the energy distribution along the chain was determined when two ions in the chain realize point contacts to heat baths at different temperatures. This configuration can be implemented by continuously laser cooling the two individual ions to different stationary temperatures. A rich behavior was predicted as a function of the cooling rate and of the position of the laser-cooled ions within the chain. Heat transport through an ion chain that undergoes the linear to zigzag transition was experimentally measured in Ref. [292]. A theoretical study reported that the onset of zigzag order gives rise to a qualitative change in transport, from ballistic to diffusive [302]. The origin of the diffusive behavior was attributed to the nonlinearity induced by coupling between the longitudinal and transverse phonons to the zigzag modes. Moreover, the heat conductivity was found to be maximimal near the critical point, where the transverse phonon spectrum becomes gapless. Transport in a chain embedding two mass impurity defects was theoretically analyzed in Ref. [120, 341]. Here, it is argued that for certain mass ratios, the motion of the mass defects can become entangled via the resonant exchange of vibrational excitations with the rest of the chain. Heat transport in the presence of a topological defect was investigated in [351] and is reviewed in Sec. III.3. The properties of transport using the radial modes of a linear chain are at the basis of proposals for quantum information with trapped ion chains [323].
@@
In microtraps, where the ion structure is designed by the external potential, transport experiments realized the coherent exchange of phononic excitations [40, 144, 143]. This experimental progress demonstrates the capability of controlling and tailoring phonon hopping between the ions and sets the basis for simulating many-body dynamics of strongly-correlated bosons [282] as well as for revealing quantum features of transport [24]. Further work has been devoted to transport controlled by the coupling with the internal degrees of freedom of the ions. We here mention the realization of vibrationally-assisted transfer of energy [136] and the simulation of an effective Jaynes-Cummings type of dynamics [320] in order to achieve effects such as phonon-blockade and designed phonon-phonon interactions [76, 264].
@@ III.3 Topological defects and metastability
Figure 43: (Color online) Experimentally observed kinks in an ion chain of about 30 ions. The kinks (red ellipses) are created during a radial quench of a crystal and imaged with exposure times of 40 ms. (a) Uniform zigzag and localized kink for trap aspect ratio $\nu/\nu_{t}\approx 8$. (b) Zigzag and extended kink for $\nu/\nu_{t}\approx 5.5$. Adapted from 270.
As mentioned in Sec. II.3.2, manifestations of prethermalization in ion Coulomb crystals such as defects and dislocations can be imaged on a CCD camera. They are typically created by means of temporal transformations of the trap potential, or by rapidly quenching the temperature while laser cooling the ion cloud to a crystal. In low dimensional crystals, and especially in the zigzag chain, it is possible to experimentally monitor the formation and the dynamics of individual defects. This has led to a series of insightful studies, some of which are reviewed in the rest of this Section.
Figure 43 reports some examples of defects (kinks) measured in a zigzag chain: the upper panel displays a so-called localized kink. The lower panel shows an example of extended kink, within which the ions form two parallel ion chains. Other kinds of kinks are composed by dislocations in all three dimensions [235, 261]. In [235] the formation of kinks was observed by laser cooling a cloud of 24Mg+ ions in a linear Paul trap into a zigzag configuration. By imaging the ion positions, kinks were measured at the center of the zigzag chain with an occurrence rate that depended on the number of ions. The kinks decayed by propagating from the center to the edges of the zigzag chain on a timescale of order 10 s, exceeding the oscillation period of the axial trap by five orders of magnitude.
@@
Defects such as those shown in Fig. 43 were also observed in experiments where the trap aspect ratio was slowly varied across the linear-zigzag transition while the chain was continuously laser cooled [110, 270, 289, 355]. They were reproduced by molecular dynamics simulations in the pseudopotential approximation, showing that micromotion is not essential in determining their onset and metastability (see Sec. II.3.2 for comparison). Their metastable nature identifies them as topological defects. Interestingly, distant localized kinks interact, as verified by a comparison between experimental data and molecular dynamics simulations [202].
@@
A classification of kinks was carried out in [202] using tools of topological degree theory. The stability of a given defect is determined by the analysis of the eigenvalues of the corresponding Hessian matrix; a bifurcation occurs every time an eigenvalue of the Hessian changes sign. Figure 44 illustrates the resulting metastable and unstable equilibrium configurations of one planar kink as a function of the trap aspect ratio, for a fixed number of ions [202].
@@
The dynamics of a kink is characterized by localized modes, which include a translational mode as well as internal modes. Theoretical analysis of the internal modes of both localized and extended kinks predict well-resolved spectroscopic features and long coherence times, even at Doppler-cooling temperatures [200]. The translational energy of the topological defects was studied by calculating the Peierls-Nabarro potential, namely, the potential energy of the soliton as a function of the kink center position [202, 270]. Numerical simulations of energy transport in ion chains with a topological defect reported energy localization at the kink [350].
@@
Figure 44: Schematic illustration of bifurcations and stability of one-kink configurations in a planar crystal of 31 ions. The possible ion configurations are depicted as a function of the ratio of confinement in the axial and one of the transverse directions, where confinement in the second transverse direction is assumed to be very strong. Solid lines indicate stable configurations (local minima of the potential), dashed lines are unstable solutions. The arrows illustrate the spatial oscillations of the normal mode which crosses zero energy at the bifurcation and generates the bifurcating solutions. From 202.
The spectroscopic features of the localized defects were measured in [42] by parametric heating. Experimental observations show features that suggest directed transport as a function of the topological charge [42]. The stability of the extended kink as a function of the trap ratio and its dynamics were extensively analyzed in [188], unveiling typical features of static friction, see Sec. III.5.
The prospect of controlling the formation and dynamics of topological defects in ion chains has inspired several theoretical proposals. In Ref. [201], it is argued that discrete solitons in a chain confined to a ring geometry can generate entanglement between ions. In a linear Paul trap, when the edge ions are heat point contacts at two different temperatures, the presence of a kink can substantially modify the heat flow, giving rise to a pronounced temperature gradient at its location [349]. Together, these insights suggest that the interactions between kinks and normal modes of the chain can modify and even tailor the transport properties (see also Sec. III.5).
Mass impurities embedded in the chain give rise to an additional class of topological defects [202]. It has been argued that the combination of mass impurities and static electric fields can allow for deterministic manipulation and creation of kinks [202]. Theoretical studies further predict that interactions between mass impurities and the surrounding chain can generate entanglement between the localized modes associated with these defects [120]. Chains containing mass or charge impurities are routinely realized in the ion-trap laboratories [116, 183, 149]. Structural phase transitions in the presence of mass impurities have been characterised in [301]. Dual species ion crystals can be deterministically prepared [376]. By positioning impurities at selected locations within the chain, the dispersion spectrum can be designed [335], with applications in quantum logic [252, 155] and metrology [317, 155].
We conclude by mentioning structural defects realized by combining spin-dependent forces and individual ion excitations in ion Coulomb crystals [74, 12, 211, 116, 220]. Such dynamics can give rise to entanglement between vibrational and internal degrees of freedom [74, 12, 134] and employed for quantum technological applications [134, 223].
@@ III.4 Slow quenches and Kibble-Zurek mechanism
Experimental imaging of kinks in ion Coulomb chains enables a statistical analysis of their formation as a function of the experimental parameters [110, 235, 270, 289, 355]. On this basis, the authors of [75, 51] argued that one can test the so-called Kibble-Zurek (KZ) hypothesis by performing slow ramps across the linear-zigzag structural transition.
The KZ hypothesis was first formulated in an attempt to describe inhomogeneities at the cosmological scale as defects formed during the slow expansion of the universe [185] and then proposed as a paradigm for adiabatic expansion across a continuous, symmetry breaking phase transition [378]. It predicts the scaling of the excess heat produced by a slow ramp of a control field across a second-order phase transition, thereby connecting the equilibrium critical exponents with features of the out-of-equilibrium dynamics.
In the experiments, the statistics of defect formation was measured after slowly ramping the trap aspect ratio across the symmetry-breaking linear-zigzag transition [110, 289, 355]. The behavior of the excess heat as a function of the ramp velocity appeared to follow power-law behavior [110, 289, 355], compatible with the KZ hypothesis. In this Section we review the theoretical model and the experimental results.
In [75] the dynamics of defect formation was modeled starting from the Landau model of Sec. II.1.4 for the transverse displacement field $\psi$ and included the dissipative dynamics of laser cooling by means of a damping and a Langevin force (see [288] for a related semiclassical model). For a chain aligned along the $x$ axis and assuming the motion along $z$ is frozen out, the displacement field is along the $y$ direction and its dynamics is governed by the equation
@@
On the right-hand side, $\eta$ is the damping rate and $\varepsilon(t)$ is the Langevin force (recall: $\langle\varepsilon(t)\rangle=0$ and $\langle\varepsilon(t)\varepsilon(t^{\prime})\rangle=(2\eta k_{\mathrm{B}}T/m)\delta(t-t^{\prime})$ where $T$ is the stationary temperature of Doppler cooling [337, 246]). The equation of motion on the left-hand side was obtained by taking the continuous limit of Eq. (45) with $V_{0}$ given by Eq. (43). Here, the first two terms describe a wave-like propagation with velocity $v(x)$, which is position dependent when the chain’s charge distribution is inhomogeneous (see e.g. Eq. (38)). The parameter $\delta(x,t)\approx\nu_{t}(t)^{2}-\nu_{t}^{(c)}(x)^{2}$ gives the amplitude of the local force in the harmonic approximation, where $\nu_{t}^{(c)}(x)$ is a position-dependent critical trap frequency, at which the chain becomes locally unstable  [^4] For a chain in a linear Paul trap, at the center the functional $\nu_{t}^{(c)}(x)$ is maximum and takes the value $\nu_{t}^{(c)}(0)=\nu_{t}^{(c)}$ (see Sec. II.1.4), while it decreases with the distance $|x|$ from the chain’s center.. The term $2\mathcal{A}\psi^{3}$ is the force in next order, see Eq. (43).
Figure 45: (Color online) Density of defects $d$ for an ion chain in a linear trap as a function of the inverse of the sweeping rate (a) in the overdamped and (b) in the underdamped regime. The circles are obtained from a molecular dynamics simulation with $50$ ions in a linear Paul trap, the defects are only counted in the central 30 ions. The red solid line is the fitting function $d\sim\tau_{Q}^{-\alpha}$ giving (a) $\alpha=0.995$ and (b) $\alpha=1.427$. Both results are consistent with the predictions of the inhomogeneous KZ mechanism. The curves saturate mostly because of finite size effects. From 50.
@@
In a linear Paul trap the ion density along $x$ is non-uniform. For long chains there is a propagating front, from the center to the edges, with which $\nu_{t}(x)$ crosses the local critical value. This front tends to give rise to independent spatial regions where the transition occurs, and thus to nucleation of defects. The speed of the propagating front should be compared with the velocity $v(x)$, at which a perturbation propagates from the center to the edges of the chain causally connecting the chain. Both quantities depend on the local density but with different functional behaviors. This leads to different scaling exponents, namely, $\alpha=1$ for the overdamped regime and $\alpha=4/3$ for the underdamped dynamics [75, 50]. These scalings are confirmed by molecular dynamics simulations, as visible in Fig. 45.
@@
Figure 46: (Color online) Defect formation rates in a chain of 16 40Ca+ ions by changing the axial trap frequency across the linear-zigzag transition. Double-logarithmic plot of the measured density of defects $d$ versus the rate of change of $\omega_{\rm ax}$ at the critical point, where $\omega_{\rm ax}$ is the trap confinement frequency along the direction of the ion chain. All circles correspond to 60000 measurements at a radial trap anisotropy of $1.03$. The fitted function of the form $d\propto 1/\tau_{Q}^{\alpha}$ (red solid line) gives an exponent $\alpha=2.68\pm 0.06$. The constant offset visible at lower ramping rates stems from background gas collisions. The saturation is due to the maximum number of defects being limited by the system size. For comparison, the rate of defects measured with a higher radial trap anisotropy of 1.05 is plotted (squares), showing a loss of defects but a similar fitted exponent of $\alpha=2.62\pm 0.15$ (dotted line). Solid data points are used for the fits. The blue stars depict the result of molecular dynamics simulations. Adapted from 355
@@
Experimentally the slow quench was performed in three different platforms, with different ion numbers and species. The trap aspect ratio was varied across the linear-zigzag transition by either sweeping the transverse [289] or the axial frequency [355], or both simultaneously [110]. In all experiments the ions were initially cooled to the Doppler limit and the cooling laser was on during the whole duration of the quench. The density of defects was measured after the quench using a CCD camera per sweep and averaging over several sweeps, yielding consistent results. Figure 46 shows the resulting statistics of kinks [355]. The solid and dashed lines show the function $d\sim\tau_{Q}^{-8/3}$, which fits the interval in which the sweep rate exceeds the rate of background collisions but is still sufficiently small so that defect generation does not saturate. This scaling is compatible with the measurements reported in [110] and was also reported in the defect statistics measured in [289] on a larger parameter interval. Molecular dynamics simulations accounting for the experimental setup reproduced the results of Ref. [75, 50] for other ramps than the experimental ones and reported the exponent $\alpha=8/3$ for the experimental regime [110, 289]. This power-law scaling has been attributed to a form of KZ mechanism, even though a theoretical model justifying this scaling has not been yet identified, see [51].